# Approaching Stable Quark Matter

### v1

@@ 7 Discussion and conclusions
Parallel to the LQCD studies at finite density, LQCD simulations at finite temperature are another way to potentially measure the bag parameter and determine the stability of quark matter together with pQCD calculations. For instance, Ref. [30] has reviewed the lattice data from Refs. [31, 32, 33] and performed a fit on the trace anomaly $\epsilon-3p$ using the $B$ parameter and the leading order pQCD contribution to obtain $B^{1/4}\simeq 200\leavevmode\nobreak\ {\rm MeV}$. One could extend this study by adopting higher-order pQCD-calculated results and treating $X$ as a model parameter, similar to what we have done in this study, to infer the stability of quark matter. More recently, Refs. [34, 35, 36] have performed simulations at even higher temperatures and analyzed the corresponding trace anomalies, which can be used to perform a more robust fit on $B$. By combining the LQCD studies at finite density and finite temperature, a more precise understanding of the bag parameter and thus the stability of quark matter could potentially be obtained.
In conclusion, we have first classified two possible relations between the hadron and quark phases of the baryon matter, which can be used to classify the existence scenarios of ordinary neutron stars, quark stars, and hybrid stars. By combining the pQCD+CS calculations and LQCD simulations for isospin-dense matter, we have inferred the sensitivity on the bag parameter $B$ and used it to constrain the possible existence of stable 2+1-flavor and/or 2-flavor quark matter. Promoting the ratio of the renormalization scale over the weighted sum of the quark chemical potentials to a model parameter $X$, the current LQCD calculations for isospin-dense matter are used to constrain the 2-dimensional parameter space of $X$ and $B^{1/4}$. The 90% CL region for $X$ and $B^{1/4}$ overlaps with the pQCD-trustable region, which hints that the quark matter stability issue could be addressed by the LQCD+pQCD framework. Furthermore, an upper bound on $B^{1/4}\lesssim 160$ MeV has been obtained, which is just around the corner from addressing the stability of quark matter. For comparison, we have also worked out the more general thermodynamic bounds on the quark matter energy-per-baryon and $B$ and found that they provide a weaker but complementary constraint.
@@ Acknowledgments
$\alpha_{s}(\mu_{R})$ $=$ $\frac{4\pi}{\beta_{0}L}\left(1-\frac{2\beta_{1}}{\beta_{0}^{2}}\frac{\ln L}{L}\right),\leavevmode\nobreak\ \qquad\mbox{with}\quad L=\ln\left(\mu_{R}^{2}/\Lambda_{\overline{\rm MS}}^{2}\right)\leavevmode\nobreak\ ,$ (A.1)
$m(\mu_{R})$ $=$ $m(2\,{\rm GeV})\left(\frac{\alpha_{s}(\mu_{R})}{\alpha_{s}(2\,{\rm GeV})}\right)^{\gamma_{0}/\beta_{0}}\times\frac{1+A_{1}\frac{\alpha_{s}(\mu_{R})}{\pi}+\frac{A_{1}^{2}+A_{2}}{2}\left(\frac{\alpha_{s}(\mu_{R})}{\pi}\right)^{2}}{1+A_{1}\frac{\alpha_{s}(2\,{\rm GeV})}{\pi}+\frac{A_{1}^{2}+A_{2}}{2}\left(\frac{\alpha_{s}(2\,{\rm GeV})}{\pi}\right)^{2}}\leavevmode\nobreak\ ,$ (A.2)
@@
$A_{1}=-\frac{\beta_{1}\gamma_{0}}{2\beta_{0}^{2}}+\frac{\gamma_{1}}{4\beta_{0}},\qquad A_{2}=\frac{\gamma_{0}}{4\beta_{0}^{2}}\left(\frac{\beta_{1}^{2}}{\beta_{0}}-\beta_{2}\right)-\frac{\beta_{1}\gamma_{1}}{8\beta_{0}^{2}}+\frac{\gamma_{2}}{16\beta_{0}}\leavevmode\nobreak\ .$ (A.3)
@@
$d_{A}=N_{c}^{2}-1,\leavevmode\nobreak\ \qquad C_{A}=N_{c},\leavevmode\nobreak\ \qquad C_{F}=\frac{N_{c}^{2}-1}{2N_{c}}\leavevmode\nobreak\ ,$
$\beta_{0}=\frac{11C_{A}-2N_{f}}{3},\qquad\qquad\beta_{1}=\frac{17}{3}C_{A}^{2}-C_{F}N_{f}-\frac{5}{3}C_{A}N_{f}\leavevmode\nobreak\ ,$
$\beta_{2}=\frac{2857}{216}C_{A}^{3}+\frac{1}{4}C_{F}^{2}N_{f}-\frac{205}{72}C_{A}C_{F}N_{f}-\frac{1415}{216}C_{A}^{2}N_{f}+\frac{11}{36}C_{F}N_{f}^{2}+\frac{79}{216}C_{A}N_{f}^{2}\leavevmode\nobreak\ ,$
$\gamma_{0}=3C_{F},\qquad\qquad\qquad\gamma_{1}=C_{F}\left(\frac{97}{6}C_{A}+\frac{3}{2}C_{F}-\frac{5}{3}N_{f}\right)\leavevmode\nobreak\ ,$
@@
$\hskip 42.67912pt+\,C_{A}N_{f}\left[-\frac{278}{27}-24\zeta(3)\right]-\frac{35}{27}N_{f}^{2}\Bigg{\}}\leavevmode\nobreak\ .$ (A.4)
@@
$u=\sqrt{\mu^{2}-m^{2}}\,,\leavevmode\nobreak\ \hat{u}=u/\mu\,,\leavevmode\nobreak\ \hat{m}=m/\mu\,,\leavevmode\nobreak\ \hat{v}=1-\hat{m}\,,\leavevmode\nobreak\ z=\hat{u}-\hat{m}^{2}\ln\left(\frac{1+\hat{u}}{\hat{m}}\right)\leavevmode\nobreak\ .$ (A.5)
@@
$\hskip 14.22636pt+d_{A}\left[\frac{4}{3}\left(N_{f}-\frac{11C_{A}}{2}\right)\ln\frac{\mu_{R}}{2\mu_{i}}-\frac{142}{9}C_{A}+\frac{17}{2}C_{F}+\frac{22}{9}N_{f}\right]\frac{g^{4}(\mu_{R})}{(4\pi)^{4}}\Bigg{\}}\leavevmode\nobreak\ ,$
@@
$\frac{\Omega^{m}}{V}$ $=\mathcal{M}_{1}+\frac{g^{2}(\mu_{R})}{(4\pi)^{2}}\mathcal{M}_{2}+\frac{g^{4}(\mu_{R})}{(4\pi)^{4}}\mathcal{M}_{3}\leavevmode\nobreak\ ,$ (A.7)
$\mathcal{M}_{1}$ $=-\frac{N_{c}\,\mu^{4}}{24\pi^{2}}\left(2\hat{u}^{3}-3z\hat{m}^{2}\right)\leavevmode\nobreak\ ,$ (A.8)
$\mathcal{M}_{2}$ $=-\frac{d_{A}\,\mu^{4}}{4\pi^{2}}\left(-6z\hat{m}^{2}\ln\frac{\mu_{R}}{m}+2\hat{u}^{4}-4z\hat{m}^{2}-3z^{2}\right)\leavevmode\nobreak\ ,$ (A.9)
@@
$\hskip 14.22636pt-N_{f}\left(4\hat{u}^{4}-13z\hat{m}^{2}-6z^{2}\right)\Bigg{]}\ln\frac{\mu_{R}}{m}+C_{A}\left(-\frac{11}{3}\ln\frac{\hat{m}}{2}-\frac{71}{9}+\mathcal{G}_{1}(\hat{m})\right)$
$\hskip 14.22636pt+C_{F}\left(\frac{17}{4}+\mathcal{G}_{2}(\hat{m})\right)+N_{f}\left(\frac{2}{3}\ln\frac{\hat{m}}{2}+\frac{11}{9}+\mathcal{G}_{3}(\hat{m})\right)+\mathcal{G}_{4}(\hat{m})\Bigg{\}}\leavevmode\nobreak\ ,$ (A.10)
$\mathcal{G}_{1}(\hat{m})$ $=32\pi^{4}\hat{m}^{2}\Big{[}-0.01863+0.02038\hat{m}^{2}-0.03900\hat{m}^{2}\ln\hat{m}$
$\hskip 14.22636pt+0.02581\hat{m}^{2}\ln^{2}\hat{m}-0.03153\hat{m}^{2}\ln^{3}\hat{m}+0.01151\hat{m}^{2}\ln^{4}\hat{m}\Big{]}\leavevmode\nobreak\ ,$ (A.11)
$\mathcal{G}_{2}(\hat{m})$ $=32\pi^{4}\hat{m}^{2}\Big{[}-0.1998-0.04797\ln\hat{m}+0.1988\hat{m}^{2}-0.3569\hat{m}^{2}\ln\hat{m}$
$\hskip 14.22636pt+0.3043\hat{m}^{2}\ln^{2}\hat{m}-0.1611\hat{m}^{2}\ln^{3}\hat{m}+0.09791\hat{m}^{2}\ln^{4}\hat{m}\Big{]}\leavevmode\nobreak\ ,$ (A.12)
$\mathcal{G}_{3}(\hat{m})$ $=32\pi^{4}\hat{m}^{2}\Big{[}-0.05741-0.02679\ln\hat{m}-0.002828\ln^{2}\hat{m}+0.05716\hat{m}^{2}$
$\hskip 14.22636pt-0.08777\hat{m}^{2}\ln\hat{m}+0.0666\hat{m}^{2}\ln^{2}\hat{m}-0.02381\hat{m}^{2}\ln^{3}\hat{m}+0.01384\hat{m}^{2}\ln^{4}\hat{m}\Big{]}\leavevmode\nobreak\ ,$ (A.13)
$\mathcal{G}_{4}(\hat{m})$ $=32\pi^{4}\hat{m}^{2}\bigg{[}0.07823+0.0388\ln\hat{m}+0.004873\ln^{2}\hat{m}-0.07822\hat{m}^{2}+0.1183\hat{m}^{2}\ln\hat{m}$
$\hskip 14.22636pt-0.08755\hat{m}^{2}\ln^{2}\hat{m}+0.03293\hat{m}^{2}\ln^{3}\hat{m}-0.01644\hat{m}^{2}\ln^{4}\hat{m}\bigg{]}\leavevmode\nobreak\ ,$ (A.14)
@@
$\frac{\Omega^{x}_{\rm VM}}{V}$ $=d_{A}\frac{g^{4}(\mu_{R})}{(4\pi)^{2}}\frac{m^{4}}{12\pi^{4}}\sum_{i=1}^{N_{l}}I_{x}\left(\frac{\mu_{i}}{m+\mu_{i}}\right)\leavevmode\nobreak\ ,$ (A.15)
@@
$\hskip 14.22636pt+\frac{\ln(1-t)}{t(1-t)^{2}}\left[1.005-0.272t(1-t)+0.154t(1-t)^{2}\right]\Bigg{\}}\leavevmode\nobreak\ ,$ (A.16)
@@
$\hskip 14.22636pt+\mu^{4}\Bigg{[}I_{13}\left(2\ln\left(\frac{g}{4\pi}\right)-\frac{1}{2}\right)+\frac{1}{2}\left(I_{20}+\frac{16}{3}(1-\ln 2)\ln 2I_{21}+I_{22}\left(\hat{m},\vec{\hat{\mu}}^{2}\right)\right)\Bigg{]}\Bigg{\}}\leavevmode\nobreak\ ,$
@@
$\Omega=\Omega^{m=0}+\Omega^{m}+\Omega^{x}_{\rm VM}+\Omega_{\rm ring}\leavevmode\nobreak\ .$ (A.18)
@@
$\Omega_{\rm massless}=\Omega^{m=0}+\Omega_{\rm ring}(m=0)\leavevmode\nobreak\ .$ (A.19)
@@
2-flavor $:\leavevmode\nobreak\ \mu_{u}=\mu_{d}-\mu_{e}\,,\leavevmode\nobreak\ \frac{2}{3}n_{u}-\frac{1}{3}n_{d}-n_{e}=0\leavevmode\nobreak\ ,$ (A.20)
2+1-flavor $:\leavevmode\nobreak\ \mu_{d}=\mu_{s}\,,\leavevmode\nobreak\ \mu_{u}=\mu_{d}-\mu_{e}\,,\leavevmode\nobreak\ \frac{2}{3}n_{u}-\frac{1}{3}n_{d}-\frac{1}{3}n_{s}-n_{e}=0\leavevmode\nobreak\ ,$ (A.21)
@@
$p_{\rm pQCD}(\mu,X)=-\frac{1}{V}\left[\Omega(\mu,X)-\Omega(\mu=\mu_{0},X)\right]\leavevmode\nobreak\ .$ (A.22)
Before concluding this section, we remark that one can in principle also consider the scenarios without the $\beta$-equilibrium requirement or electrons and thus charge neutrality is satisfied solely by the different quark compositions. As pointed out in Ref. [10], electrons usually play a minor role in the overall thermodynamic properties of the system. For illustration, we show the $[p(\mu_{\text{B}})]^{1/4}$ curves of the 2-flavor $X=1.6,2.2$ scenarios with and without electrons in Figure 8, from which one can see that the two scenarios produce similar predictions.
Figure 8: The $[p(\mu_{\text{B}})]^{1/4}$ curves of the 2-flavor $X=1.6,2.2$ scenarios with and without electrons. For the scenarios without electrons, charge neutrality is satisfied solely by the $u$- and $d$-quark compositions or $2n_{u}=n_{d}$.

### v3

@@ 7 Discussion and conclusions
Parallel to the LQCD studies at finite density, LQCD simulations at finite temperature are another way to potentially measure the bag parameter and determine the stability of quark matter together with pQCD calculations. For instance, Ref. [46] has reviewed the lattice data from Refs. [47, 48, 49] and performed a fit on the trace anomaly $\epsilon-3p$ using the $B$ parameter and the leading order pQCD contribution to obtain $B^{1/4}\simeq 200\penalty 10000\ {\rm MeV}$. One could extend this study by adopting higher-order pQCD-calculated results and treating $X$ as a model parameter, similar to what we have done in this study, to infer the stability of quark matter. More recently, Refs. [50, 51, 52] have performed simulations at even higher temperatures and analyzed the corresponding trace anomalies, which can be used to perform a more robust fit on $B$. By combining the LQCD studies at finite density and finite temperature, a more precise understanding of the bag parameter and thus the stability of quark matter could potentially be obtained.
In conclusion, we have first classified two possible relations between the hadron and quark phases of the baryon matter, which can be used to classify the existence scenarios of ordinary neutron stars, quark stars, and hybrid stars. By combining the pQCD+CS calculations and LQCD simulations for isospin-dense matter, we have inferred the sensitivity on the bag parameter $B$ and used it to constrain the possible existence of stable 2+1-flavor and/or 2-flavor quark matter. Promoting the ratio of the renormalization scale over the weighted sum of the quark chemical potentials to a model parameter $X$, the current LQCD calculations for isospin-dense matter are used to constrain the 2-dimensional parameter space of $X$ and $B^{1/4}$. The 90% CL region for $X$ and $B^{1/4}$ overlaps with the pQCD-trustable region, which hints that the quark matter stability issue could be addressed by the LQCD+pQCD framework. Furthermore, an upper bound on $B^{1/4}\lesssim 160$ MeV has been obtained, which is just around the corner from addressing the stability of quark matter. On the other hand, the $\langle m_{q}\overline{q}q\rangle$ condensates provide a minimum contribution to or a lower bound on $B$, which excludes the possible existence of the stable 2+1-flavor quark matter while leaving a tight parameter space for the stable 2-flavor quark matter up to two caveats: the still unknown $\mathcal{O}(\alpha_{s}^{3})$ (even higher-order) corrections to $p_{\rm pQCD}$ and the possible partial restoration of the quark and gluon condensates. For comparison, we have also worked out the more general thermodynamic bounds on the quark matter energy-per-baryon and $B$ and found that they provide a weaker but complementary constraint.
@@ Acknowledgments
$\alpha_{s}(\mu_{R})$ $=$ $\frac{4\pi}{\beta_{0}L}\left(1-\frac{2\beta_{1}}{\beta_{0}^{2}}\frac{\ln L}{L}\right),\penalty 10000\ \qquad\mbox{with}\quad L=\ln\left(\mu_{R}^{2}/\Lambda_{\overline{\rm MS}}^{2}\right)\penalty 10000\ ,$ (A.1)
$m(\mu_{R})$ $=$ $m(2\,{\rm GeV})\left(\frac{\alpha_{s}(\mu_{R})}{\alpha_{s}(2\,{\rm GeV})}\right)^{\gamma_{0}/\beta_{0}}\times\frac{1+A_{1}\frac{\alpha_{s}(\mu_{R})}{\pi}+\frac{A_{1}^{2}+A_{2}}{2}\left(\frac{\alpha_{s}(\mu_{R})}{\pi}\right)^{2}}{1+A_{1}\frac{\alpha_{s}(2\,{\rm GeV})}{\pi}+\frac{A_{1}^{2}+A_{2}}{2}\left(\frac{\alpha_{s}(2\,{\rm GeV})}{\pi}\right)^{2}}\penalty 10000\ ,$ (A.2)
@@
$A_{1}=-\frac{\beta_{1}\gamma_{0}}{2\beta_{0}^{2}}+\frac{\gamma_{1}}{4\beta_{0}},\qquad A_{2}=\frac{\gamma_{0}}{4\beta_{0}^{2}}\left(\frac{\beta_{1}^{2}}{\beta_{0}}-\beta_{2}\right)-\frac{\beta_{1}\gamma_{1}}{8\beta_{0}^{2}}+\frac{\gamma_{2}}{16\beta_{0}}\penalty 10000\ .$ (A.3)
@@
$d_{A}=N_{c}^{2}-1,\penalty 10000\ \qquad C_{A}=N_{c},\penalty 10000\ \qquad C_{F}=\frac{N_{c}^{2}-1}{2N_{c}}\penalty 10000\ ,$
$\beta_{0}=\frac{11C_{A}-2N_{f}}{3},\qquad\qquad\beta_{1}=\frac{17}{3}C_{A}^{2}-C_{F}N_{f}-\frac{5}{3}C_{A}N_{f}\penalty 10000\ ,$
$\beta_{2}=\frac{2857}{216}C_{A}^{3}+\frac{1}{4}C_{F}^{2}N_{f}-\frac{205}{72}C_{A}C_{F}N_{f}-\frac{1415}{216}C_{A}^{2}N_{f}+\frac{11}{36}C_{F}N_{f}^{2}+\frac{79}{216}C_{A}N_{f}^{2}\penalty 10000\ ,$
$\gamma_{0}=3C_{F},\qquad\qquad\qquad\gamma_{1}=C_{F}\left(\frac{97}{6}C_{A}+\frac{3}{2}C_{F}-\frac{5}{3}N_{f}\right)\penalty 10000\ ,$
@@
$\hskip 42.67912pt+\,C_{A}N_{f}\left[-\frac{278}{27}-24\zeta(3)\right]-\frac{35}{27}N_{f}^{2}\Bigg\}\penalty 10000\ .$ (A.4)
@@
[…]
@@
$\hskip 14.22636pt+d_{A}\left[\frac{4}{3}\left(N_{f}-\frac{11C_{A}}{2}\right)\ln\frac{\mu_{R}}{2\mu_{i}}-\frac{142}{9}C_{A}+\frac{17}{2}C_{F}+\frac{22}{9}N_{f}\right]\frac{g^{4}(\mu_{R})}{(4\pi)^{4}}\Bigg\}\penalty 10000\ ,$
@@
$\frac{\Omega^{m}}{V}$ $=\mathcal{M}_{1}+\frac{g^{2}(\mu_{R})}{(4\pi)^{2}}\mathcal{M}_{2}+\frac{g^{4}(\mu_{R})}{(4\pi)^{4}}\mathcal{M}_{3}\penalty 10000\ ,$ (A.7)
$\mathcal{M}_{1}$ $=-\frac{N_{c}\,\mu^{4}}{24\pi^{2}}\left(2\hat{u}^{3}-3z\hat{m}^{2}\right)\penalty 10000\ ,$ (A.8)
$\mathcal{M}_{2}$ $=-\frac{d_{A}\,\mu^{4}}{4\pi^{2}}\left(-6z\hat{m}^{2}\ln\frac{\mu_{R}}{m}+2\hat{u}^{4}-4z\hat{m}^{2}-3z^{2}\right)\penalty 10000\ ,$ (A.9)
@@
[…]
@@
$\frac{\Omega^{x}_{\rm VM}}{V}$ $=d_{A}\frac{g^{4}(\mu_{R})}{(4\pi)^{2}}\frac{m^{4}}{12\pi^{4}}\sum_{i=1}^{N_{l}}I_{x}\left(\frac{\mu_{i}}{m+\mu_{i}}\right)\penalty 10000\ ,$ (A.15)
@@
$\hskip 14.22636pt+\frac{\ln(1-t)}{t(1-t)^{2}}\left[1.005-0.272t(1-t)+0.154t(1-t)^{2}\right]\Bigg\}\penalty 10000\ ,$ (A.16)
@@
$\hskip 14.22636pt+\mu^{4}\Bigg[I_{13}\left(2\ln\left(\frac{g}{4\pi}\right)-\frac{1}{2}\right)+\frac{1}{2}\left(I_{20}+\frac{16}{3}(1-\ln 2)\ln 2I_{21}+I_{22}\left(\hat{m},\vec{\hat{\mu}}^{2}\right)\right)\Bigg]\Bigg\}\penalty 10000\ ,$
@@
$\Omega=\Omega^{m=0}+\Omega^{m}+\Omega^{x}_{\rm VM}+\Omega_{\rm ring}\penalty 10000\ .$ (A.18)
@@
$\Omega_{\rm massless}=\Omega^{m=0}+\Omega_{\rm ring}(m=0)\penalty 10000\ .$ (A.19)
@@
2-flavor $:\penalty 10000\ \mu_{u}=\mu_{d}-\mu_{e}\,,\penalty 10000\ \frac{2}{3}n_{u}-\frac{1}{3}n_{d}-n_{e}=0\penalty 10000\ ,$ (A.20)
2+1-flavor $:\penalty 10000\ \mu_{d}=\mu_{s}\,,\penalty 10000\ \mu_{u}=\mu_{d}-\mu_{e}\,,\penalty 10000\ \frac{2}{3}n_{u}-\frac{1}{3}n_{d}-\frac{1}{3}n_{s}-n_{e}=0\penalty 10000\ ,$ (A.21)
@@
$p_{\rm pQCD}(\mu,X)=-\frac{1}{V}\left[\Omega(\mu,X)-\Omega(\mu=\mu_{0},X)\right]\penalty 10000\ .$ (A.22)
Figure 8: The ratios of the $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{1}]$ ($\Omega^{(1)}$) and $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{2}]$ ($\Omega^{(2)}$) contributions to the leading-order term $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{0}]$ ($\Omega^{(0)}$) in Eq. (A.6) for $X=2$ and neglecting quark masses.
To investigate the perturbativity of the pQCD calculations, we assume a simple massless three-flavor scenario (in which $\mu_{\text{B}}=3\mu$) and show in Figure 8 the ratios of the $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{1}]$ ($\Omega^{(1)}$) and $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{2}]$ ($\Omega^{(2)}$) terms to the $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{0}]$ term ($\Omega^{(0)}$) in Eq. (A.6) for $X=2$ (and, as mentioned in Section 3, $(\frac{\alpha}{\pi})<1$ throughout the $\mu_{\text{B}}$ range of our interest), from which one can see that the individual ratios are both justifiably small. One might question about the behavior of the $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{3}]$ correction, which, if in the extreme limit takes up the positive sign with a magnitude comparable to the $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{2}]$ correction, can possibly reverse our conclusion about the stability of the 2+1-flavor quark matter (see Section 5). To answer this question, we perform a similar analysis up to $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{1}]$ and show two similar plots to Figure 6 in Figure 9, from which one can see that the 2+1-flavor quark matter is still well excluded if we take the (more reliable) lattice result for the $s$-quark condensate, while for the sum rule average value, the conclusion still holds up to $X\sim 2.2$, which already exceeds the usual maximum value of the renormalization scale parameter $X=2$ used in the literature. Although we cannot completely exclude the possibility of a quark matter system taking up $X>2$ or that the $\mathcal{O}[(\frac{\alpha_{s}}{\pi})^{3}]$ correction being large enough to invalidate the analysis of this study, we think they are unlikely to happen according to the partial results presented in Ref. [12] (which also shows the increasing trend of $\mu_{0}$ at larger $X$ values that will further solidify the exclusion of stable 2+1-flavor quark matter) and can only be answered in future studies.
Figure 9: Similar plots to Figure 6 but with pQCD only up to $\mathcal{O}(\alpha_{s}^{1})$.
Before concluding this section, we remark that one can in principle also consider the scenarios without the $\beta$-equilibrium requirement or electrons and thus charge neutrality is satisfied solely by the different quark compositions. As pointed out in Ref. [11], electrons usually play a minor role in the overall thermodynamic properties of the system. For illustration, we show the $p(\mu_{\text{B}})]^{1/4}$ curves of the 2-flavor $X=1.6,2.2$ scenarios with and without electrons in Figure 10, from which one can see that the two scenarios produce similar predictions.
Figure 10: The $p(\mu_{\text{B}})]^{1/4}$ curves of the 2-flavor $X=1.6,2.2$ scenarios with and without electrons. For the scenarios without electrons, charge neutrality is satisfied solely by the $u$- and $d$-quark compositions or $2n_{u}=n_{d}$.